OpenStudy (anonymous):

Find the exact solution, using common logarithms, and a two-decimal-place approximation of each solution. 3^(x+1)=2^(1-3x)

OpenStudy (anonymous):

\[\ln 3^{x+1} = \ln 2^{1-3x}\] \[(x+1)ln 3 = (1-3x)ln 2\] \[xln3+ln3=ln2-3xln2\] \[xln3+3xln2=ln2-ln3\] \[x(ln3+3ln2)=ln2-lnl3\] \[x=\frac{\ln2+\ln3}{3\ln2+\ln3}\]

OpenStudy (anonymous):

I guess I didn't think of it that way. Thanks for the help!

OpenStudy (anonymous):

Oh, that should be ln2-ln3 in the numerator.

OpenStudy (anonymous):

Yes, I made a typo in the last step.